OpenStudy (anonymous):

f(x)=(4x+2)^-1. Find f'(x) using the definition of the derivative

OpenStudy (anonymous):

f(x)=(4x+2)^-1 f '(x)=[-1(4x+2)^-2]4 =-4/(4x+2)^2 ans...........

OpenStudy (anonymous):

The definition of a derivative says: \[f'(x)=\lim_{\phi \rightarrow 0}\frac{(4(x+\phi)+2)^{-1}-(4x+2)^{-1}}{\phi}\] Manipulate it until you can do it by hand. I won't show it here.